OpenStudy (anonymous):

Let f(x) = x+1/x+m. Find the inverse f^-1(x). For what value of m is f(x) = f^-1(x)?

OpenStudy (anonymous):

To get f(x)^-1, flip the numerator and denominator around f(x)^-1=(x+m)/(x+1)

OpenStudy (anonymous):

i got y(x-1) = 1-xm is that right @M0j0jojo

OpenStudy (anonymous):

?

OpenStudy (freckles):

To find the inverse of f(x)=y you need to solve f(x)=y for x.

OpenStudy (freckles):

\[y=\frac{x+1}{x+m} \text{ solve for x}\]

OpenStudy (freckles):

I will help you out with the first couple of steps ... Multiply (x+m) on both sides (x+m)y=x+1 xy+my=x+1 xy-x=1-my

OpenStudy (freckles):

See if you can figure out what comes next... Remember you are trying to solve for x.

OpenStudy (freckles):

Actually based on what you said you know the next step x(y-1)=1-my now divide both sides by (y-1)

OpenStudy (freckles):

\[x=\frac{1-my}{y-1} \\ f^{-1}(y)=\frac{1-my}{y-1} \\ \text{ So } f^{-1}(x)=\frac{1-mx}{x-1}\]

OpenStudy (freckles):

So you want to find m such that \[\frac{1-mx}{x-1}=\frac{x+1}{x+m}\]

OpenStudy (freckles):

If you just compare the bottoms what do you think m needs to be?

OpenStudy (anonymous):

is the answer -1?

OpenStudy (freckles):

yes if m=-1 on bottom we have x-1 on both bottom's and we also see that works on top because 1+x is x+1

OpenStudy (anonymous):

ok thank you soooo much :)